# ReconX: Reconstruct Any Scene from Sparse Views with Video Diffusion Model

### v2

@@ 5.2 Comparison with Feed-Forward based Baselines
Table 3: Quantitative comparisons with per-scene optimization based methods on MipNeRF 360  and Tank and Temples, and DL3DV. We evaluate the reconstruction performance with different input views for each scene.
5.3 Comparison with Per-Scene Optimization based Baselines
@@
To verify the capability of ReconX in sparse-view reconstruction in more challenging outdoor settings, we compare with per-scene optimization-based methods in different input views (i.e., , 2, 3, 6, and 9 views) in Table 3 and more visual comparisons in Figure 4. We observe that our method outperforms all the other per-scene optimization baselines in PSNR, SSIM, and LPIPS scores. As shown in Figure 4, we find that the baselines produce extremely blurry results in only two view settings with noisy camera estimations. In contrast, by unleashing the generative power of the video diffusion model, our ReconX can create more observations from only two sparse views and ensures high-quality novel view rendering, avoiding local minima issues. To further demonstrate our superiority, we compare in 3-view setting even with recent works CAT3D (Gao et al., 2024) and ReconFusion (Wu et al., 2024b) that incorporate generative prior to mitigate ill-posed sparse view reconstruction in Table 5 and Figure 5. Since the codes for CAT3D and ReconFusion are not available, we downloaded the results directly from the project pages using three input views as provided in their papers. The results show that ReconX can produce higher-frequency details in novel views.
Figure 4: Qualitative comparison with per-scene optimization based methods on Mip-Nerf 360 and Tank and Temples. With two sparse views as input, our ReconX achieves much better reconstruction quality compared with baselines.
Figure 5: More qualitative comparisons with three sparse views as input. (a) comparison with DNGaussian and SparseNeRF with random selected input views. (b) comparison with CAT3D and ReconFusion on input views selected with heuristic loss (Wu et al., 2024b).
5.4 Ablation Study and Analysis
We carry out ablation studies on RealEstate10K to analyze the design of our ReconX framework in Table 4 and Figure 6. A naive combination of pretrained video diffusion model and Gaussian Splatting is regarded as the “base”. Specifically, we ablate on the following aspects of our method: 3D structure condition, DUSt3R initialization, confidence-aware optimization, and LPIPS loss. The results indicate that the omission of any of these elements leads to a degradation in terms of quality and consistency. Notably, the basic combination of original video diffusion model and 3DGS leads to significant distortion of the scene. The absence of 3D structure condition causes inconsistent generated frames especially in distant input views, resulting in blur and artifact issues. The lack of confidence-aware optimization leads to suboptimal results in some local detail areas. Adding LPIPS loss in confidence-aware 3DGS optimization would provide clearer rendering views. Moreover, we ablate the impact of DUSt3R and video diffusion priors in Figure 7. Although the point cloud may not include enough high-quality information, such coarse 3D structure is sufficient to guide the video diffusion in our ReconX to fill in the distortions, occlusions or missing regions. This demonstrates that our ReconX has learned a comprehensive understanding of the 3D scene and can generate high-quality novel views from imperfect conditional information and exhibit robustness to the point cloud conditions. This illustrates the effectiveness of our overall framework (Figure 2), which drives generalizable and high-fidelity 3D reconstruction given only sparse views as input.
@@ 5.4 Ablation Study and Analysis
Table 4: Quantitative results of ablation study. We report the quantitative metrics in ablations of our framework in real-world data (Zhou et al., 2018).
6 Conclusion
In this paper, we introduce ReconX, a novel sparse-view 3D reconstruction framework that reformulates the inherently ambiguous reconstruction problem as a generation problem. The key to our success is that we unleash the strong prior of video diffusion models to create more plausible observations frames for sparse-view reconstruction. Grounded by the empirical study and theoretical analysis, we propose to incorporate 3D structure guidance into the video diffusion process for better 3D consistent video frames generation. What’s more, we propose a 3D confidence-aware scheme to optimize the final 3DGS from generated frames, which effectively addresses the uncertainty issue. Extensive experiments demonstrate the superiority of our ReconX over the latest state-of-the-art methods in terms of high quality and strong generalizability in unseen data.
Limitations and Future Work. Although ReconX achieves remarkable reconstruction results in novel viewpoints, the quality still seems to be limited by the backbone as we choose the U-Net based diffusion model DynamiCrafter (Xing et al., 2023). We expect that this issue can be solved with open-sourced larger video diffusion models (e.g., DiT-based framework). In the future, it is interesting to integrate 3DGS optimization directly with the video generation model, enabling more efficient end-to-end 3D scene reconstruction. We are also interested in exploring consistent 4D scene reconstruction. We believe that ReconX provides a promising research direction to craft intricate 3D worlds from video diffusion models and hope it will inspire more works in the future.
Appendix A Preliminaries
Video Diffusion Models. Diffusion models (Ho et al., 2020; Song et al., 2020) have emerged as the cutting-edge paradigm to generate high-quality videos. These models learn the underlying data distribution by adding and removing noise on the clean data. The forward process aims to transform a clean data sample $\bm{x}_{0}\sim p(\bm{x})$ to a pure Gaussian noise $\bm{x}_{T}\sim\mathcal{N}(0,I)$, following the process:
$\bm{x}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\quad\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{1}),$ (8)
where $\bm{x}_{t}$ and $\bar{\alpha}_{t}$ denotes the noisy data and noise strength at the timestep $t$. The denoising neural network $\epsilon_{\theta}$ is trained to predict the noises added in the forward process, which is achieved by the MSE loss:
$\mathcal{L}=\mathbb{E}_{\bm{x}\sim p,\epsilon\sim\mathcal{N}(0,I),c,t}\left[\left\|\epsilon-\epsilon_{\theta}\left(\bm{x}_{t},t,c\right)\right\|_{2}^{2}\right],$ (9)
where $c$ represents the embeddings of conditions like text or image prompt. For the video diffusion models, Latent Diffusion Models (LDMs) (Rombach et al., 2022), which compress images into the latent space, are commonly employed to mitigate the computation complexity while maintaining competitive performance.
3D Gaussian Splatting. 3DGS (Kerbl et al., 2023) represents a scene explicitly by utilizing a set of 3D Gaussian spheres, achieving a fast and high-quality rendering. A 3D Gaussian is modeled by a position vector $\bm{\mu}\in\mathbb{R}^{3}$, a covariance matrix $\bm{\Sigma}\in\mathbb{R}^{3\times 3}$, an opacity $\alpha\in\mathbb{R}$, and spherical harmonics (SH) coefficient $\bm{c}\in\mathbb{R}^{k}$ (Ramamoorthi & Hanrahan, 2001). Moreover, the Gaussian distribution is formulated as the following:
$G(x)=e^{-\frac{1}{2}(x-\bm{\mu})^{T}\bm{\Sigma}^{-1}(x-\bm{\mu)}},$ (10)
where $\bm{\Sigma}=\bm{R}\bm{S}\bm{S}^{T}\bm{R}^{T}$, $\bm{S}$ denotes the scaling matrix and $\bm{R}$ is the rotation matrix.
In the rendering stage, the 3D Gaussian spheres are transformed into 2D camera planes through rasterization (Zwicker et al., 2001). Specifically, given the perspective transformation matrix $\bm{W}$ and Jacobin of the projection matrix $\bm{J}$, the 2D covariance matrix in the camera space is computed as
$\bm{\Sigma}^{{}^{\prime}}=\bm{J}\bm{W}\bm{\Sigma}\bm{W}^{T}\bm{J}^{T}.$ (11)
End-to-end Dense Unconstrained Stereo. DUSt3R (Wang et al., 2024a) is a new model to predict a dense and accurate 3D scene representation solely from image pairs without any prior information about the scene. Given two unposed images $\{\bm{I}_{1},\bm{I}_{2}\}$, this end-to-end model is trained to estimate the point maps $\{P_{1,1},P_{2,1}\}$ and confidence maps $\{\mathcal{C}_{1,1},\mathcal{C}_{2,1}\}$, which can be utilized to recover the camera parameters and dense point cloud. The training procedure for view $v\in\{1,2\}$ is formulated as a regression loss:
Appendix B More Implementation Details
Implementation of PosEmb. The PosEmb implemented in our paper is a column-wise positional embedding function: $\mathbb{R}^{3}\rightarrow\mathbb{R}^{C}$, where $C$ is the dimension of embedding. More specifically, the PosEmb function is implemented as follows: (1) Fixed Sinusoidal Basis: The basis $\mathbf{e}$ is a 3D sinusoidal encoding: $\mathbf{e}=[\sin(2^{0}\pi p),\sin(2^{1}\pi p),\dots]$, where $p\in\mathbb{R}^{3}$ is the position. (2) Embedding Calculation: The input $\mathbf{x}$ is projected onto $\mathbf{e}$ and its sine and cosine are concatenated: $\mathbf{embeddings}=\text{concat}(\sin(\mathbf{proj}),\cos(\mathbf{proj}))$. (3) Learnable Transformation: The positional encoding is passed through an MLP along with the input $\mathbf{x}$: $\mathbf{y}=\text{MLP}(\text{concat}(\mathbf{embeddings},\mathbf{x}))$. In short, PosEmb combines a fixed sinusoidal encoding with a learnable MLP transformation.
More details of transformer-based encoder. For the transformer-based encoder, we encode the DUSt3R point cloud data to a fixed-length sparse representation of the point cloud. Specifically, we first employ a subsampling based on farthest point sampling (FPS) to reduce the point cloud to a smaller set of key points while retaining its overall structural characteristics. Then, we apply cross-attention between the embeddings of the original point cloud and downsampled point cloud. This mechanism can be interpreted as a form of partial self attention, where the downsampled points act as query anchors that aggregate information from the original point cloud. The encoder is not initialized from any pretrained models. Instead, it is trained jointly with the video diffusion model in an end-to-end manner. This design choice ensures that the encoder is specifically adapted to the characteristics of DUSt3R point clouds in our experiment datasets.
Camera alignment. We consider that the camera pose from DUSt3R is not aligned with the COLMAP cameras. Since we adopt the point cloud and camera poses from DUSt3R in our experiment, it is necessary for us to unify the training and testing images into the same DUSt3R coordinate system. Specifically, we process the training and testing images together through DUSt3R to obtain the corresponding camera poses, and utilize only the point maps from the training set as the initial point cloud for optimizing 3DGS.
Test view selection. In comparison with feed-forward based methods, we follow MVSplat (Chen et al., 2024a) and pixelSplat (Charatan et al., 2024) to choose test views in Easy Set. For Hard Set, we choose the frame intervals much larger (i.e., $>$ 200 frames) than Easy Set. For Tank-and-Templates and DL3DV datasets, we select the training views evenly from all the frames and use every 8th of the remaining frames for evaluation. For nine scenes in Mip-NeRF 360 dataset, we manually choose a training 9-view split of views that are uniformly distributed around the hemisphere and pointed toward the central object of interest. Then we further choose the 6- and 3-view splits to be subsets of the 9-view split.
Appendix C Theoretical Proof
Proposition 1. Let $\bm{\theta}^{*},\psi^{*}=g^{*}$ be the optimal solution of the solely image-based conditional diffusion scheme and $\bm{\tilde{\theta}}^{*},\tilde{\psi}^{*}=\{{g}^{*},\mathcal{F}^{*}\}$ be the optimal solution of diffusion scheme with native 3D prior. Suppose the divergence $\mathcal{D}$ is convex and the embedding function space $\Psi$ includes all measurable functions, we have $\mathcal{D}(q\left(\bm{x}\right)\|p_{\bm{\tilde{\theta}}^{*},\tilde{\psi}^{*}}(\bm{x}))<\mathcal{D}\left(q\left(\bm{x}\right)\|p_{\bm{\theta}^{*},\psi^{*}}(\bm{x})\right)$.
$\mathcal{D}\left(q(\bm{x})\|p_{\bm{\tilde{\theta}}^{*},\tilde{\psi}^{*}}(\bm{x})\right)$ $=\mathcal{D}\left(\mathbb{E}_{q(s)}q(\bm{x}|s)\|\mathbb{E}_{q(s)}p_{\bm{\tilde{\theta}}^{*},\tilde{\psi}^{*}}(\bm{x}|s)\right)$ (15)
$\leq\mathbb{E}_{q(s)}\mathcal{D}\left(q(\bm{x}|s)\|p_{\bm{\tilde{\theta}}^{*},\tilde{\psi}^{*}}(\bm{x}|s)\right)$
$=\mathbb{E}_{q(s)}\mathcal{D}\left(q(\bm{x}|s)\|p_{\bm{\tilde{\theta}}^{*},{g}^{*},\mathcal{F}^{*}}(\bm{x}|s)\right),$
where we incorporate an intermediate variable $s$, which represents a specific scene. $q(\bm{x}|s)$ indicates the conditional distribution of rendering data $\bm{x}$ given the specific scene $s$. According to the definition of $\bm{\tilde{\theta}}^{*},{g}^{*},\mathcal{F}^{*}$, we have:
$\mathbb{E}_{q(s)}\mathcal{D}\left(q(\bm{x}|s)\|p_{\bm{\tilde{\theta}}^{*},{g}^{*},\mathcal{F}^{*}}(\bm{x}|s)\right)$ $=\min_{\bm{{\theta}},{g},\mathcal{F}}\mathbb{E}_{q(s)}\mathcal{D}\left(q(\bm{x}|s)\|p_{\bm{{\theta}},{g},\mathcal{F}}(\bm{x}|s)\right)$ (16)
$=\min_{\bm{{\theta}}}\mathbb{E}_{q(s)}\min_{{g}(s),\mathcal{F}(s)}\mathcal{D}\left(q(\bm{x}|s)\|p_{\bm{{\theta}},{g}(s),\mathcal{F}(s)}(\bm{x})\right)$
$=\min_{\bm{{\theta}}}\mathbb{E}_{q(s)}\min_{{g},E}\mathcal{D}\left(q(\bm{x}|s)\|p_{\bm{{\theta}},{g},E}(\bm{x})\right),$
$\mathcal{D}\left(q(\bm{x})\|p_{\bm{\tilde{\theta}}^{*},\tilde{\psi}^{*}}(\bm{x})\right)$ $\leq\min_{\bm{{\theta}}}\mathbb{E}_{q(s)}\min_{{g},E}\mathcal{D}\left(q(\bm{x}|s)\|p_{\bm{{\theta}},{g},E}(\bm{x})\right)$ (17)
$<\min_{\bm{\theta},g,E}\mathcal{D}\left(q(\bm{x})\|p_{\bm{\theta},g,E}(\bm{x})\right)=\min_{\bm{\theta},g,E(\varnothing)}\mathcal{D}\left(q(\bm{x})\|p_{\bm{\theta},g,E(\varnothing)}(\bm{x})\right)$
$=\min_{\bm{\theta},\psi}\mathcal{D}\left(q(\bm{x})\|p_{\bm{\theta},\psi}(\bm{x})\right)=\mathcal{D}\left(q\left(\bm{x}\right)\|p_{\bm{\theta}^{*},\psi^{*}}(\bm{x})\right).$
The second inequality holds because given general real-world scene $s$ in any parameter $\bm{\theta}\in\Theta$, approximating $q(\bm{x}|s)$ is simpler than $q(\bm{x})$ by only tuning the encoder $E$ of $p_{\bm{\theta},g,E}$ [^1] A simple verifiable case is to optimize the parameters of 3DGS by only 2D images (solely image-based conditional learning) or using a SFM initialization from collected images (native 3D conditional learning) before optimization. The latter provides a more constrained and optimal solution space., i.e., $\min_{E}\mathcal{D}\left(q(\bm{x}|s)\|p_{\bm{{\theta}},{g},E}(\bm{x})\right)<\min_{E}\mathcal{D}\left(q(\bm{x})\|p_{\bm{\theta},g,E}(\bm{x})\right)$ holds almost everywhere (a.e.), representing $\mathcal{P}_{q(s)}\left\{\min_{E}\mathcal{D}\left(q(\bm{x}\mid s)\|p_{\bm{\theta},g,E}(\bm{x})\right)<\min_{E}\mathcal{D}\left(q(\bm{x})\|p_{\bm{\theta},g,E}(\bm{x})\right)\right\}=1$
Consequently, the proof of Proposition 1 has been done.
Appendix D More Results and Analysis
Evaluation of our extrapolation ability. As we use a pair of input views in our method, it is worthy to note that if the angular difference between the two views is too large, it is hard to ensure that the entire interpolated region falls within the visible perspective of the input views, which requires the extrapolation ability. We have evaluated it in our generalizable experiments with DTU dataset. For instance, in the case of DTU in Figure 3, we cannot see the roof area from the input views, while our ReconX is able to extrapolate and generate the red and yellow roof with 3D structure-guided generative prior. To further demonstrate the extrapolation capability of our method, we conduct a specific experiment in Figure 8. This experiment selects two views with large angular spans and highlights the extrapolated regions in the red boxes in the novel-rendered views. This emphasizes our model’s generative power to extrapolate unseen regions and extend beyond the visible input views.
More visual results in outdoor scenes. Regarding the DL3DV dataset, we trained our model on this to demonstrate its performance on outdoor scenes. Due to the limitations of feed-forward methods on this dataset, we did not present quantitative results in the main paper, as these methods fail on it. However, to highlight our model’s strengths in outdoor environments, we have included visual results in the supplementary video and have added comparisons with per-scene optimization methods in Table 3. We have also provided more visual results on DL3DV in Figure 12. We also compare our ReconX in 3D Gaussians with frame-by-frame results in Figure 10.
@@ Appendix D More Results and Analysis
More quantitative comparisons. As the data is open-sourced in ReconFusion (Wu et al., 2024b) we conduct an additional quantitative experiment in comparison with ZipNeRF (Barron et al., 2023), ZeroNVS (Sargent et al., 2023), CAT3D (Gao et al., 2024), and ReconFusion (Wu et al., 2024b). It is worth noting that the data split used in CAT3D (Gao et al., 2024) follows a heuristic loss (Gao et al., 2024) to encourage reasonable camera spacing and coverage of the central object. We observe that our ReconX is better than all baselines in Table 5.
Method 3-view 6-view 9-view
[…]
Zip-NeRF 12.77 0.271 0.705 13.61 0.284 0.663 14.30 0.312 0.633
ZeroNVS 14.44 0.316 0.680 15.51 0.337 0.663 15.99 0.350 0.655
ReconFusion 15.50 0.358 0.585 16.93 0.401 0.544 18.19 0.432 0.511
CAT3D 16.62 0.377 0.515 17.72 0.425 0.482 18.67 0.460 0.460
ReconX (Ours) 17.16 0.435 0.407 19.20 0.473 0.378 20.13 0.482 0.356
Table 5: Quantitative comparisons with more per-scene optimization based methods on MipNeRF 360. We evaluate the reconstruction performance with different input views for each scene.
Figure 10: Rendering comparison with Gaussian-based methods frame by frame.
More extrapolation ability discussion. In the main paper, we focus on unleashing the video diffusion model to generate 3D consistent views through two-view interpolation. Specifically, we have conducted experiments to verify the capability of ReconX to recover 3D scenes from large angle variance in input views (see Figure 3 and Figure 8), showing its ability to extrapolate the occlusions and correct the inaccurate geometry details given the coarse 3D structure. As the position of our conditional images in ReconX is inherently flexible, allowing us to unleash more extrapolation capability by adjusting the placement of the conditional images.

### v3

@@ V-A Experiment Setup
Figure 5: Rendering comparison of sparse-view 3D scene reconstruction with Gaussian-based methods frame by frame.
@@
TABLE IV: Quantitative comparisons with more multi-views reconstruction methods on MipNeRF 360. We evaluate the reconstruction performance with different input views for each scene.
Method 3-view 6-view 9-view
[…]
Zip-NeRF 12.77 0.271 0.705 13.61 0.284 0.663 14.30 0.312 0.633
ZeroNVS 14.44 0.316 0.680 15.51 0.337 0.663 15.99 0.350 0.655
ReconFusion 15.50 0.358 0.585 16.93 0.401 0.544 18.19 0.432 0.511
CAT3D 16.62 0.377 0.515 17.72 0.425 0.482 18.67 0.460 0.460
ReconX (Ours) 17.16 0.435 0.407 19.20 0.473 0.378 20.13 0.482 0.356
Figure 6: Qualitative results of our ReconX on outdoor scenes from DL3DV [61].
Datasets. The video diffusion model of ReconX is trained on three datasets: RealEstate-10K [62], ACID [63], and DL3DV-10K [61] based on the pretrained model. RealEstate-10K is a dataset downloaded from YouTube, which is split into 67,477 training scenes and 7,289 test scenes. The ACID dataset consists of natural landscape scenes, with 11,075 training scenes and 1,972 testing scenes. DL3DV-10K is a large-scale outdoor dataset containing 10,510 videos with consistent capture standards. For each scene video, we randomly sample 32 contiguous frames with random skips and serve the first and last frames as the input for our video diffusion model. In two-views novel view synthesis experiment, we follow MVSplat [14] and pixelSplat [12] to choose test views in Easy Set. For Hard Set, we choose the frame intervals much larger (i.e., $>$ 200 frames) than Easy Set. To further validate our strong generalizability, we also directly evaluate our method on the DTU [64], NeRF-LLFF [65], and more challenging outdoor datasets Mip-NeRF 360 [66] and Tank-and-Temples dataset [67]. For DTU, NeRF-LLFF and Tank-and-Templates datasets, we select the training views evenly from all the frames and use every 8th of the remaining frames for evaluation. For nine scenes in Mip-NeRF 360 dataset, we manually choose a training 9-view split of views that are uniformly distributed around the hemisphere and pointed toward the central object of interest. Then we further choose the 6- and 3-view splits to be subsets of the 9-view split.
Baselines and Metrics. To comprehensively demonstrate our strong capability in sparse-view reconstruction, we compare our ReconX with (a) feed-forward based methods trained from 3D scenes to learn 3D prior and (b) per-scene optimization based methods with specific priors (e.g., , depth) for sparse-view reconstruction. Specifically, we compare with NeRF-based pixelNeRF [11] and MuRF [68]; Light Field based GPNR [69] and AttnRend [70]; and the recent state-of-the-art 3DGS-based pixelSplat [12] and MVSplat [14] in feed-forward based comparisons. On the other hand, we compare with SparseNeRF [71], original 3DGS [2], and DNGaussian [72] for per-scene optimization comparisons. Furthermore, we qualitatively compare our method with more recent works CAT3D [10] and ReconFusion [15] that incorporate generative power. For quantitative results, we report the standard metrics in NVS, including PSNR, SSIM [73], LPIPS [58].
V-B 3D Scene Reconstruction from Two-Views
Comparison for small angle variance in input views. For fair comparison with two-views novel view synthesis baseline methods like MuNeRF [68], pixelSplat [12], and MVSplat [14], we first compare our reconX with baseline method from sparse views with small angle variance (see Easy Set from Table II and Fig. 3). We observe that our ReconX surpasses all previous state-of-the-art models in terms of all metrics on visual quality and qualitative perception.
Comparison for large angle variance in input views. As MVSplat and pixelSplat are much better than previous baselines, we conduct thorough comparisons with them in more difficult settings. In more challenging settings (i.e., given sparse views with large angle variance), our proposed ReconX demonstrate more significant improvement than baselines, especially in unseen and generalized viewpoints (see Hard Set from Table II and Fig. 3). This clearly shows the effectiveness of ReconX in creating more consistent observations from video diffusion to mitigate the inherent ill-posed sparse-view reconstruction problem.
Cross-dataset generalization. Unleashing the strong generative power of the video diffusion model through 3D structure condition, our ReconX is inherently superior in generalizing to out-of-distribution novel scenes. To demonstrate the strong generalizability of ReconX, we conduct two cross-dataset evaluations. For a fair comparison, we train the models solely on the RealEstate10K and directly test them on two popular NVS datasets (i.e., NeRF-LLFF [65] and DTU [64]). As shown in Cross Set from Table II and Fig. 3, the competitive baseline methods MVSplat [14] and pixelSplat [12] fail to render such OOD datasets which contain different camera distributions and image appearance, leading to dramatic performance degradation. In contrast, our ReconX shows impressive generalizability and the gain is larger when the domain gap from training and test data becomes larger.
V-C 3D Scene Reconstruction from Multi-Views
@@
To verify the capability of ReconX in sparse-view (more than two views) reconstruction in more challenging outdoor settings, we compare with multi-views reconstruction methods in different input views (i.e., , 2, 3, 6, and 9 views) in Table III and more visual comparisons in Fig. 4. We observe that our method outperforms all the other per-scene optimization baselines in PSNR, SSIM, and LPIPS scores. As shown in Fig. 4, we find that the baselines produce extremely blurry results in only two view settings with noisy camera estimations. In contrast, by unleashing the generative power of the video diffusion model, our ReconX can create more observations from only two sparse views and ensures high-quality novel view rendering, avoiding local minima issues.
TABLE V: Quantitative results of ablation study. We report the quantitative metrics in ablations of our framework in real-world data [62].
@@
To further demonstrate our superiority, we compare in even with recent works like CAT3D [10] and ReconFusion [15] that incorporate generative prior to mitigate ill-posed sparse view reconstruction. As the data is open-sourced in ReconFusion [15], we conduct an additional quantitative experiment in comparison with ZipNeRF [74], ZeroNVS [42], CAT3D [10], and ReconFusion [15]. It is worth noting that the data split used in CAT3D [10] follows a heuristic loss [10] to encourage reasonable camera spacing and coverage of the central object. We observe that our ReconX is better than all baselines in Table IV.
Regarding the DL3DV dataset, we trained our model on this to demonstrate its performance on outdoor scenes. Due to the limitations of feed-forward methods on this dataset, we did not present quantitative results in the main paper, as these methods fail on it. However, to highlight our model’s strengths in outdoor environments, we have included visual results in the supplementary video and have added comparisons with per-scene optimization methods in Table III. We have also provided more visual results on DL3DV in Fig. 6. We also compare our ReconX in 3D Gaussians with frame-by-frame results in Fig. 5.
V-D Evaluation of Extrapolation Ability
As we use a pair of input views in our method, it is worthy to note that if the angular difference between the two views is too large, it is hard to ensure that the entire interpolated region falls within the visible perspective of the input views, which requires the extrapolation ability. We have evaluated it in our generalizable experiments with DTU dataset. For instance, in the case of DTU in Fig. 3, we cannot see the roof area from the input views, while our ReconX is able to extrapolate and generate the red and yellow roof with 3D structure-guided generative prior. To further demonstrate the extrapolation capability of our method, we conduct a specific experiment in Fig. 7. This experiment selects two views with large angular spans and highlights the extrapolated regions in the red boxes in the novel-rendered views. This emphasizes our model’s generative power to extrapolate unseen regions and extend beyond the visible input views.